LHC 2012 - latest news

Week 46: Machine coordinators: Gianluigi Arduini, Jan Uythoven

Goal of this week: Luminosity production.

Date

Summary

Link to report from morning meeting

Mon 19 Nov
  • 03:11 RF problem solved.
  • 05:38 Stable beams fill 3300. Initial lumi 7.5e33 cm-2s-1.
08:30 meeting 
Sun 18 Nov
  • 00:03 Beam dump due to instability and losses on two bunches leading to BPM LSS6 trigger. ~10 pb-1 in 25 mins. Revert to smaller tune split (-0.003)
  • 02:30 Stable beams 3298. Initial luminosity 7.5e33 cm-2s-1
  • 05:45 Beam dumped by OP on request of ALICE for an urgent intervention on the cooling of their TPC.
  • End of fill 3298 producing ~70 pb-1 in 3:17 h.
  • 06:20 Access starting
  • 08:30 Access finished. Restarting
  • 10:53 STABLE BEAMS #3299. Initial luminosity 7.3e33 cm-2s-1
  • 22:37 Operator dump, 174 pb-1 in 11h44m.
  • 23:07 RF trip M1B2. Access needed.
09:00 meeting
Sat 17 Nov
  • 00:18 Sector 56 tripped while setting to stand-by after access. Problem with RQ10.R5, QPS problem suspected.
  • 02:30 Problem RQ10.R5 identified to be due to power converter problem.
  • 04:50 Access for power converter RQ10.R5.
  • 05:29 Access finished.
  • 06:42 Injecting fill 3296, clean without steering.
  • 07:35 Start ramp.
  • 08:23 Stable beams 3296, initial lumi around 7.1e33 cm-2s-1.
  • 15:30 No beam from SPS Beam dump switch to be replaced)
  • 20:40 Beam available from SPS
  • 20:50 OP dump: ~180 pb-1 in 12.5 hours.
  • 23:39 Stable beams 3297. Initial luminosity 7.3e33 cm-2s-1. Larger tune split during the squeeze (-0.005)
08:30 meeting 
Fri 16 Nov
  • 00:47: Beam loss spike at Q6R5 close to Roman Pot (running sum 1.2 s). Beam dump at 1.2 TeV.
  • Reduced bunch intensity to ~1.6x1011 p and longitudinal blow-up to 1.4 ns
  • 02:58 Beam dump during adjust: UFO like event of Roman Pot Q6L5
  • 05:16 STABLE BEAMS fill 3292. 6.3x1033 cm-2s-1
  • 08:56 Beam dump (SW interlock TCDQ reading - old data). 68 pb-1 in 3h40m.
  • 11:27 Stable beams fill 3293. Initial lumi 6.85e33 cm-2s-1. Longitudinal blow-up 1.4 ns B2 only.
  • 12:44 Beam dump, QPS trigger on RQT13.L5B2. Integrated lumi 28 pb-1 in 1h17m.
  • 14:17 Start ramp fill 3294.
  • 14:27 Dump during ramp at 2.86 GeV,  dump due to slow losses at Q6.R5.
  • 17:14 Start ramp fill 3295.
  • 18:29 Operator dump before colliding in LHCb. Steering problems.
  • 20:43 Injection fill 3296. Stop, problem corrector RCBH.23L6B1.
  • 21:00 Access for RCBH.23L6B1 and warm up beam screen Q6.R5 near TOTEM.
08:30 meeting
Thu 15 Nov
  • 04:30 OP dump. End of fill 3287. 137 pb-1 in 9h01m.
  • No transfer line steering. Profit from reduced PS extraction kicker pulse length.
  • 06:58 Stable beams fill 3288. Initial luminosity ~6.6e33 cm-2s-1.
  • 2fb-1 delivered to LHCb
  • 18:47-20:00 TOTEM test with separated beams in CMS
  • 20:02 OP dump 176 (ATLAS) / 170 (CMS) pb-1 in 13 h. More than 21 fb-1 delivered to ATLAS/CMS
  • 22:53 Beam loss spike at Q6R5 close to Roman Pot (running sum 5.6 s). Beam dump at 2.8 TeV.
08:30 meeting
Wed 14 Nov
  • 03:30 Stable beams fill 3286, initial lumi's up7.5 e33 cm-2s-1.
    ALICE at 6.6e30 cm-2s-1
  • 13:52 Beam dump when moving TOTEM Roman Pot in to 30 mm setting.
    162 pb-1 in 10h20m.
  • 14:45 No beam fron LINAC2.
  • 16:13 Problem with PS water cooling.
  • 17:07 Injectors back. Optimising beams in injectors. Steering.
  • 19:31 Stable beams fill 3287, initial lumi's around 6.4e33 cm-2s-1.
    ALICE 5.6e30 cm-2s-1
08:30 meeting
Tue 13 Nov
  • 06:30 Cryogenics back pre-cycle.
  • 08:00 Injecting pilots, steering.  
  • 09:00 No beam from injectors, problem with Booster distributer.
  • 11:35 Booster back. Check injection 144 bunches: OK.
  • 12:00 Set-up of gated BBQ for B2.
  • 12:30 Continue filling for physics: very clean.
  • 13:00 Start ramp fill 3285.
  • 13:56 Stable beams, initial lumi's around 7.2 e33, ALICE 3.4e30 cm-2s-1.
  • 24:00 End of fill 3285. OP dump.153 pb-1 in ~10h10m.
08:30 meeting
Mon 12 Nov
  • 04:50 Beam dump, trip of RF line 4B, integrated L ~120 pb-1.
  • Access for RF.
  • 07:00 Injecting. FBCT phasing. Line steering.
  • 09:59 Beam dunp LHCb on injection. Reduce satellites in PS.
  • 10:30 Transvers damper module H2B2 down and RF lines tripped.
  • 13:24 ADT problem fixed. Injecting for physics.
  • 13:52 Start ramp.
  • 13:55 Cryo problem in point 4. Cryo ok will be lost in few minutes.
  • 14:05 Beam dumped due to cryogenics point 4.
  • Cryo recovery.
08:30 meeting